# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v3

@@ 2 Series and generating functions in $\mathbb{C}$
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then the power series $\sum_{n\geq 0}u_{n}x^{n}$ absolutely converges for every $x\in\mathbb{C}$ with $|x|<1$.
In fact, for the absolute convergence it suffices (and is necessary) by Hadamard’s formula that $\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1$.
Definition 2.14 ($U(1)_{1}=+\infty$)
@@
$\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}=+\infty\,,}$
we say that $U(1)_{1}=+\infty$.
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $V(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-x}$, then
$\lim_{n\to\infty}u_{n}=V(1)_{1}\,.$
Proof. Since $\frac{1}{1-x}=1+x+x^{2}+\dots$, we have $u_{n}=v_{0}+v_{1}+\dots+v_{n}$. $\Box$
Quantities $U(1)_{1}$ are handled by means of the classical theorem due to N. Abel, see [1] and [28, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
Theorem 2.16 (Abel’s 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists then the function
$F_{U}(x)=\sum_{n\geq 0}u_{n}x^{n}\colon\{x\in\mathbb{C}\colon\;|x|<1\}\to\mathbb{C}\,$,}
see Proposition 2.13, has the limit
$\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=U(1)_{1}\,.$
The following folklore result is in fact another variant of Abel’s theorem, incomparable with Theorem 2.16. We omit the easy proof.
Theorem 2.17 (Abel’s 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$. If $U(x)$ absolutely converges for every $x\in[0,1)$, then the equality
$\lim_{x\to 1}F_{U}(x)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}\ \ (\in[0,\,+\infty)\cup\{+\infty\})$}
always holds.
Compared to Theorem 2.16, the new case is when both limits equal $\infty$.
@@
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then $U(1)_{1}$ exists and
@@
By the limit transition $n\to\infty$ we get that that $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Iterating this result we easily generalize it to $k$-term linear combinations.
@@
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$, $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions,
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$
(see proofs of Propositions 2.25 and 2.27 for details of similar arguments). Theorem 2.16 and properties of limits of functions give
$U(1)_{1}$ $=$ $\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
@@
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $V(0)\neq 0$. If $U(x)=\frac{1}{V(x)}$ and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.19 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and division by $V(1)_{1}$ yields the stated formula. $\Box$
@@
Let $U(x),V(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0)\neq 0$. If $\frac{1}{V(x)}, U(x)U(x)$ and $V(x)$ absolutely converge for $x\in[0,1)$, and if $V(1)_{1}=+\infty$, then
$U(1)_{1}=0\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. By Propositions 2.4 and 2.5, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$. Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$. Two applications of Theorem 2.17 and properties of limits of functions give
$U(1)_{1}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)=\lim_{x\to 1}\frac{1}{F_{V}(x)}=\frac{1}{\lim_{x\to 1}F_{V}(x)}}$
$=$ ${\textstyle\frac{1}{\lim_{n\to\infty\sum_{j=0}^{n}v_{j}}}=\frac{1}{+\infty}=0\,.}$
$\Box$
Proposition 2.22
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$ and $W(0)\neq 0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.18, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Division by $W(1)_{1}$ yields the stated formula. $\Box$
Proposition 2.23
Let $U(x),V(x),W(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0),W(0)>0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(x)$, $V(x)$ and $W(x)$ absolutely converge for $x\in[0,1)$, if $F_{V}(x)$ is bounded on $[0,1)$ and if $W(1)_{1}=+\infty$, then $U(1)_{1}$ exists and
$U(1)_{1}=1\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Using the assumptions and Propositions 2.2, 2.4 and 2.6 we get that $F_{U}(x)^{2}\cdot F_{W}(x)=F_{W}(x)-F_{V}(x)$ for every $x\in[0,\,1)$. Since $F_{W}(0)>0$ and $F_{W}(x)$ is non-decreasing on $[0,1)$, we can divide by $F_{W}(x)$ and get that $F_{U}(x)^{2}=1-\frac{F_{V}(x)}{F_{W}(x)}$ for every $x\in[0,\,1)$. Note that $\lim_{x\to 1}F_{V}(x)=a>0$ because $F_{V}(0)>0$ and $F_{V}(x)$ is non-decreasing and bounded from above. Theorem 2.17 and properties of limits of functions give
$(U(1)_{1})^{2}$ $=$ ${\textstyle\big(\lim_{x\to 1}F_{U}(x)\big)^{2}=\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)}$
$=$ ${\textstyle 1-\frac{\lim_{x\to 1}F_{V}(x)}{\lim_{x\to 1}F_{W}(x)}=1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
Hence $U(1)_{1}=1$ because $U(1)_{1}\geq 0$. $\Box$
@@
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then $U(1)_{2}$ exists and
@@
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}}v_{n}$ and $\sum_{n\in\mathbb{N}}w_{n}$ and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
@@
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
Proposition 2.19 for $n=2$ and item 1 of Proposition 2.25 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to Franz (Franciszek) Mertens (1840–1927).
@@
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
@@
    $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. It is clear that $V(1)_{2}$ exists. Since $|V(1)_{2}-1|\leq c<1$, we get $V(1)_{2}\neq 0$. We consider the series
@@
We show that $S$ absolutely converges. It follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most $1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}$.
For $m\in\mathbb{N}$ we consider the series
@@
By Proposition 2.6 it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
@@
By Proposition 2.4, $S$ has also sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
2. Indeed,

### v6

@@ 2 Series and generating functions in $\mathbb{C}$
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. Then $U(x)\in H_{1}$ iff
$\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1\,.$
Thus, for example, the existence of $U(1)_{1}$ is a sufficient (but not necessary) condition for $U(x)\in H_{1}$.
Note that if $U\in\mathbb{C}[[x]]$ then $(U)_{\xi}\in H_{1}$ either for all or no $\xi\in\mathbb{C}_{1}$.
Definition 2.14 ($U(1)_{1}=\pm\infty$)
@@
$\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}=+\infty\,$,}
we say that $U(1)_{1}=+\infty$. We similarly define that $U(1)_{1}=$-$\infty$.
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$, and let $\xi\in\mathbb{C}^{*}$. If $(V)_{1/\xi}(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-\xi x}$, then
$\lim_{n\to\infty}\xi^{-n}u_{n}=(V)_{1/\xi}(1)_{1}\,.$
Proof. Let $n\in\mathbb{N}_{0}$. Since $\frac{1}{1-\xi x}=1+\xi x+\xi^{2}x^{2}+\dots$, we have
$u_{n}=v_{0}\xi^{n}+v_{1}\xi^{n-1}+\dots+v_{n}\xi^{0}=v_{0}\cdot\xi^{n}+v_{1}\xi^{-1}\cdot\xi^{n}+\dots+v_{n}\xi^{-n}\cdot\xi^{n}\,.$
Dividing by $\xi^{n}$ and taking the limit for $n\to\infty$, we get the stated formula. $\Box$
We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems, which we state only for the radius of convergence $R\geq 1$. The first theorem is classical, see [1, 2] and [29, Chapter II.7].
Theorem 2.16 (Abel 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists, so that $U(x)\in H_{1}$ by Proposition 2.13, then
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
The second Abelian theorem is rarely mentioned, but see [2] for a simple proof. It is given only for the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
Theorem 2.17 (Abel 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}[[x]]\cap H_{1}$. If $U(1)_{1}=\pm\infty$ then
$\lim_{x\nearrow 1}F_{U}(x)=\pm\infty\,$,
with equal signs.
The third Abelian theorem allows only nonnegative real coefficients. It has the advantage over the previous two theorems that it is an equivalence. Now one can go not only from the sum to the functional limit, but also in the reverse. We omit the easy proof.
Theorem 2.18 (Abel 3)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist (and are possibly $\infty$) and are equal,
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
@@
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then
@@
Taking the limit for $n\to\infty$ we get that that $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Here the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”. We will use such abbreviated formulations. This result can be easily generalized to linear combinations with more than two terms. Next we consider products of formal power series.
@@
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1$, $n\in\mathbb{N}$, formal power series. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions we have
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}_{<1}$
(see proofs of Propositions 2.24 and 2.27 for details of similar arguments). Using Theorem 2.16 and properties of limits of functions we get
$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}F_{U}(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
@@
Let $U(x),V(x)\in\mathbb{C}[[x]]$. If $U(x)=\frac{1}{V(x)}$, and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and dividing by $V(1)_{1}$ gives the stated formula. $\Box$
@@
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.19, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Dividing by $W(1)_{1}$ we get the stated formula. $\Box$
@@
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then
@@
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
@@
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}\,$,}
where the inner $\sum$ means sum, the outer $\sum$ means series, and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
For nonnegative real coefficients we can go in the reverse, from the product to the factors.
Proposition 2.25
Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]$. Suppose that $U(x)=V(x)W(x)$, that $U(x)$, $V(x)$ and $W(x)$ are in $H_{1}$, $V(x)W(x)\neq 0$ and that $U(1)_{2}$ exists. Then $V(1)_{2}$ and $W(1)_{2}$ exist, and
$U(1)_{2}=V(1)_{2}W(1)_{2}\,.$
Proof. By Theorem 2.18 and the assumption $V(x)W(x)\neq 0$ we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}F_{V}(x)>0$ and the same for $W(x)$. Since $F_{U}=F_{V}F_{W}$ on $\mathbb{C}_{<1}$, we have
$U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}F_{U}(x)$
$=$ $\lim_{x\nearrow 1}F_{V}(x)\cdot\lim_{x\nearrow 1}F_{W}(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
In the second equality we used Theorem 2.18. The third equality holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to F. Mertens.
@@
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then
@@
    The sum $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. From the assumption on $c$ it follows that $V(1)_{2}$ exists. From
$|V(1)_{2}-1|\leq c<1$
we get $V(1)_{2}\neq 0$. We consider the series
@@
We show that $S$ absolutely converges. This follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most
$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.$}
Let $m\in\mathbb{N}$. We consider the series
@@
By Proposition 2.6, it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum
$1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}\,.$}
@@
Proposition 2.4 shows that $S$ also has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
2. Indeed, the sum